A plug-in transformer supplies 9.00 V to a video game system.
(a) How many turns are in its secondary coil if its input voltage is 120 V and the primary coil has 383 turns?
turns

(b) What is its input current when its output is 1.39 A?
A
Answer #1

The ratio rule for transformers states that

\frac{Vp}{Vs} = \frac{Np}{Ns}

Here V and N denotes the voltage and number of turns. P and S denotes primary and secondary.

a)

Since Vs= 9V and Vp = 120V and Np = 383

Thus Ns = 28.725 turns.

b)

Since V = IR

thus \frac{Vp}{Vs} = \frac{Ip}{Is}

here I is the current

Therefore Is = 1.39 A

therefore Ip = 18.53 A